### v1

# Local consistency and axioms of functional dependence

Abstract
Local consistency arises in diverse areas, including Bayesian statistics, relational databases, and quantum foundations. Likewise, the notion of functional dependence arises in all of these areas. We adopt a general approach to study logical inference in a setting that enables both global inconsistency and local consistency. Our approach builds upon pairwise consistent families of $K$-relations, i.e, relations with tuples annotated with elements of some positive commutative monoid. The framework covers, e.g., families of probability distributions arising from quantum experiments and their possibilistic counterparts. As a first step, we investigate the entailment problem for functional dependencies (FDs) in this setting. Notably, the transitivity rule for FDs is no longer sound, but can be replaced by two novel axiom schemes. We provide a complete axiomatisation and a PTIME algorithm for the entailment problem of unary FDs. In addition, we explore when contextual families over the Booleans have realisations as contextual families over various monoids.
1 Introduction
Local consistency is an important concept in various areas such as Bayesian statistics, relational databases and quantum foundations. In broad terms, local consistency refers to a family of partial structures (such as marginal distributions or projections) that agree on their overlapping parts. Local consistency is usually a desirable property to enforce or at least a minimal requirement or observed phenomenon. It is often contrasted with global consistency, which demands the existence of a single underlying structure whose projections yield the given family. In practice, local consistency sometimes serves as a proxy for global consistency, especially when the latter is infeasible to check or maintain explicitly.
In Bayesian statistics, local consistency is central to the representation of global probability distributions via factorisation. For example, Bayesian networks encode a global distribution as a family of marginal and conditional distributions that must be consistent with each other. Such factorisations are very useful as they facilitate computationally feasibly probabilistic and approximate reasoning without the need to directly access or store the global distribution.
In relational databases, local consistency arises when data is spread across multiple tables with overlapping attribute sets. The motivation here is not computational speed-up—storing everything in a single table would eliminate costly joins in query evaluation—but data integrity, as decomposing data helps to prevent anomalies, such as update anomalies. Notably, consistency in relational databases is typically enforced via integrity constraints that are local in nature. E.g., data about a person may be distributed across multiple tables, which we expect to be globally consistent; yet this consistency is usually maintained via primary keys and foreign keys that act on individual tables or pairs of tables only.
In quantum theory, local consistency is a principle that places an upper bound on the outcomes appearing in physical experiments: any experimental setup must produce a family of probability distributions that satisfies local consistency. However, not all locally consistent families are compatible with quantum predictions.
This intricate relationship between local and global consistency naturally gives rise to many interesting questions. In domains where local consistency is used as a proxy for global consistency, it is natural to ask: under what conditions does local consistency entail global consistency? If local consistency is a necessary but not sufficient condition for observable phenomena, one may query about the reasons that account for this gap.
For instance, quantum theory offers remarkably accurate predictions on experimental outcomes, yet it may not provide a satisfactory explanation—in terms of fundamental principles or laws of nature —as to why certain locally consistent distributions are not physically realisable. Several efforts have been made to explain this discrepancy in terms of simple fundamental principles. A notable example is the principle of information causality (?), which uses the language of information theory to constrain admissible distributions.
In the context of databases and probability distributions, in turn, much effort has gone into characterising when local consistency guarantees global consistency. The so-called local-to-global consistency refers to the situation where the underlying structure (such as a database schema or collection of variable sets) guarantees that every locally consistent family (of relations or distributions) can be extended to a globally consistent one. The characterisations of this property are typically formulated in the language of graph theory. For relational databases it is known that a database schema satisfies the local-to-global consistency property if and only if it forms an acyclic hypergraph (?). Similar characterisations have been also shown in the context of distributions and multisets (??). More recently, these results have been generalised in the setting of $K$-relations, which are relations whose tuples are annotated with elements from some positive and commutative monoid (?). In this case, the assumption of acyclicity of the schema is not always sufficient but a more involved characterisation is required.
Our Contributions. In this article, we turn the tables and treat local consistency not as a proxy for global consistency, but as a foundational concept in its own right. Analogously to (?), we adopt the useful abstraction of $K$-relations that allows us to treat families of relations and distributions uniformly. We also redirect our focus from domains such as information theory and graph theory to logical consequence, and investigate how it is impacted by the shift from global to local consistency. As a test case, we study the logical entailment between functional dependencies (FDs). Our approach bears similarity to the logic of local inference of ? (?) and to quantum team logic of (?) which are geared for applications in quantum theory, where local consistency does not generally entail the existence of a global model (cf. Bell’s theorem).
After defining our theoretical framework, we investigate when a given family of locally consistent relations $\mathcal{R}$ can be enriched to a family of locally consistent $K$-relations (with the same support). Here, we give a complete characterisation for the case, where $K$ is cancellative and $\mathcal{R}$ is a contextual family over what we refer to as a chordless-cycle context set. Moreover, we establish a sound and complete axiomatisation for the entailment of unary FDs in the setting of locally consistent $K$-relations. In particular, we observe that the usual transitivity rule of FDs becomes unsound in this context. Interestingly, the transitivity rule disperses into an infinite collection of weaker rules that we name chain and cycle rules. In the standard unirelational context these rules are all derivable from the transitivity rule, and one instance of a chain rule derives the transitivity rule. Finally, we establish that the derivability question of our axiomatisation can be decided in polynomial time. As an instantiation of our result, we obtain insight on the entailment of FDs in settings that lack global consistency, such as in quantum foundations.
2 Preliminaries
For a natural number $n$, we write $[n]:=\{0,1,\dots,n-1\}$. We fix a countably infinite set Var of variables. Given a set $A$, an assignment (of $A$) is a function $s$ that maps a finite set $D\subseteq\textrm{Var}$ of variables to some values (in $A$). We call $D$ the domain of $s$ and write $\mathrm{Dom}(s)$. The set of all assignments $s\colon D\to A$ is denoted $\textrm{As}(D,A)$.
A monoid is an algebraic structure $K=(K,+,0)$, where $+$ is associative and 00 is the identity element of $+$. $K$ is positive if $a+b=0$ entails $a=0=b$, for all $a,b\in K$, and (left) cancellative if $a+b=a+c$ entails $b=c$. We associate each monoid $K$ with its natural order $\leq$, defined by $a\leq b\iff\exists c:a+c=b$. The natural order of a monoid is reflexive and transitive, meaning that it is a preorder. In this paper, $K$ will always denote a positive commutative non-trivial monoid (the trivial monoid contains only 00).
Given a finite set of variables $D\subseteq\textrm{Var}$ and a finite set $A$, a $K$-relation over $D$ is a function $R\colon\textrm{As}(D,A)\to K$. We write $\mathrm{Vars}(R)$ for the set of variables $D$ and $\mathrm{Dom}(R)$ for the set $\textrm{As}(D,A)$. The support of $R$ is the relation $\mathrm{Supp}(R)\coloneqq\{s\in\mathrm{Dom}(R)\mid R(s)\neq 0\}$.
For a relation $R$ and an element $c\in K$ of a monoid $K$, we denote by $cR$ the $K$-relation of support $R$ such that $(cR)(s)=c$ for all $s\in R$. For two $K$-relations with the same domain $X$, $R\colon X\to K$ and $S\colon X\to K$, we define the $K$-relation $T=R+S$ as follows: $T\colon X\to K$ and $T(s)=R(s)+S(s)$ for all $s\in X$. For a set of variables $D^{\prime}$, the $K$-relation $R\!\restriction\!D^{\prime}$ is the marginalisation of the $K$-relation $R$ to the variables in $D^{\prime}$. This means that $R\!\restriction\!D^{\prime}$ is a $K$-relation $\textrm{As}(D^{\prime},A)\to K$ s.t.
$(R\!\restriction\!D^{\prime})(s):=\sum_{\begin{subarray}{c}s^{\prime}\in\mathrm{Dom}(R)\\ s=s^{\prime}\!\restriction\!D^{\prime}\end{subarray}}R(s^{\prime}),$
where $s\!\restriction\!D^{\prime}$ is the restriction of the assignment $s$ to $D^{\prime}$ and $\sum$ is the aggregate sum of the monoid.
In addition to the general case, we consider the Boolean monoid $\mathbb{B}=(\{0,1\},\lor,0)$ and the monoids of non-negative reals $\mathbb{R}_{\geq 0}=([0,\infty),+,0)$ and natural numbers $\mathbb{N}=(\mathbb{N},+,0)$ with their usual addition.
3 $K$-relations and Contextual Families
In quantum information theory, a context is a set of attributes that can be measured together. More formally, a context $C$ is a set of variables and context set $\mathcal{C}$ is a downward closed set of contexts (i.e, $C\in\mathcal{C}$ and $C^{\prime}\subseteq C$ implies $C^{\prime}\in\mathcal{C}$). We often represent a context set $\mathcal{C}$ in terms of its maximal elements; for instance, we may write $\mathcal{C}=\{\{x,y\},\{y,z\}\}$ to denote $\mathcal{C}=\{\{x,y\},\{y,z\},\{x\},\{y\},\{z\},\emptyset\}$. To remove clutter, we often write a context $\{x_{1},\dots,x_{n}\}$ as $x_{1}\dots x_{n}$.
Let $R$ and $S$ be $K$-relations over sets of variables $D$ and $D^{\prime}$, respectively. We say that $R$ and $S$ are consistent if $R\upharpoonright(D\cap D^{\prime})=S\upharpoonright(D\cap D^{\prime})$. A family of $K$-relations $\mathcal{R}$ is locally consistent if all pairs of $K$-relations from $\mathcal{R}$ are consistent, and globally consistent if there exists a $K$-relation $R$ such that $\mathrm{Vars}(R)=\bigcup_{S\in\mathcal{R}}\mathrm{Vars}(S)$ and $R\!\restriction\!\mathrm{Vars}(S)=S$ for all $S\in\mathcal{R}$.
Definition 1 (Contextual $K$-families).
Let $\mathcal{C}$ be a context set. A contextual $K$-family over $\mathcal{C}$ is a locally consistent set of $K$-relations $\mathcal{R}$ containing exactly one $K$-relation $R$ with domain $C$, for each $C\in\mathcal{C}$. The $K$-relation of $\mathcal{R}$ with domain $C$ is referred to as $\mathcal{R}_{C}$. The set $\mathcal{C}$ is the context set of $\mathcal{R}$, denoted $\textrm{Conts}(\mathcal{R})$.
The support $\mathrm{Supp}(\mathcal{R})$ of $\mathcal{R}$ is defined as $\{\mathrm{Supp}(\mathcal{R}_{C})\mid C\in\mathcal{C}\}$. A contextual $\mathbb{B}$-family is simply called a contextual family, and identified with its support.
Proposition 2.
If $\mathcal{R}$ is a contextual $K$-family, then $\mathrm{Supp}(\mathcal{R})$ is a contextual family.
If $\mathcal{R}$ is a contextual family and $\alpha\in K$, we set $\alpha\mathcal{R}$ as $\{\alpha\mathcal{R}_{C}\mid C\in\mathcal{C}\}$. In contrast to the above, if $\mathcal{R}$ is a contextual family, $\alpha\mathcal{R}$ is not necessarily a contextual $K$-family (it may violate local consistency; cf. Ex. 5). Let $\mathcal{R}$ and $\mathcal{S}$ be contextual $K$-families over a shared context set $\mathcal{C}$. We set $\mathcal{R}+\mathcal{S}$ as $\{\mathcal{R}_{C}+\mathcal{S}_{C}\mid C\in\mathcal{C}\}$. Clearly, local consistency is preserved under this operation.
Proposition 3.
If are $\mathcal{R}$ and $\mathcal{S}$ are contextual $K$-families over a shared context set $\mathcal{C}$, then $\mathcal{R}+\mathcal{S}$ is a contextual $K$-family.
A contextual family $\mathcal{R}$ is sometimes identified with the set of its assignments $\bigcup{\mathcal{R}}=\bigcup_{C\in\mathcal{C}}\mathcal{R}_{C}$. Similarly, a contextual $K$-family $\mathcal{S}$ can be viewed as a single function #:Supp($\#:\bigcup{\mathrm{Supp}(\mathcal{S})}\to K_{\neq 0}$, where $K_{\neq 0}=\{a\in K\mid a\neq 0\}$. A contextual family is called $K$-realisable if it is the support of some contextual $K$-family.
Example 4.
[…]
Student Teacher
Alice Charlie
Bob David
Teacher Course
Charlie Math
David CS
Course Student
Math Alice
CS Alice
CS Bob
Figure 1: Teachers, courses, and students.
3.1 $K$-realisability of Contextual Families
Let us now turn our attention to the $K$-realisability of contextual families. It is natural to ask whether or not any contextual family is the support of a contextual $K$-family. This turns out not to be the case for $K=\mathbb{R}_{\geq 0}$. A counterexample—in our terminology, a contextual family over the context set $\{ab,ab^{\prime},a^{\prime}b,a^{\prime}b^{\prime}\}$—is given in (?, Proposition 9.1). Example 4 provides another counterexample, as we argue next.
Example 5.
Consider the contextual family in Fig. 1. For a contradiction, suppose this contextual family is the support of some contextual $\mathbb{R}_{\geq 0}$-family, represented as a function #$\#\colon S\to\mathbb{R}_{>0}$, where $S$ is the set of assignments in Fig. 1. By local consistency, $\pi$ extends uniquely to assignments on individual variables. Now
#$\#(\texttt{Student}\mapsto\text{Bob})=\pi(\texttt{Teacher}\mapsto\text{David})$
$=$ #$\#(\texttt{Course}\mapsto\text{CS})$
[…]
$>$ #$\#(\texttt{Student}\mapsto\text{Bob}),$
where the inequality follows since each value of ##\## is strictly positive. Thus the contextual family of Fig. 1 cannot be obtained as the support of any contextual $\mathbb{R}_{\geq 0}$-family.
Next, we completely characterise which contextual families are supports of some contextual $K$-families for cancellative monoids $K$ and context sets that form a chordless cycle (defined below). We show that a contextual family $\mathcal{R}$ is the support of some contextual $K$-family if and only if its associated overlap projection graph has an edge cycle cover; both terms are defined below.
For what follows, we treat addition and subtraction of natural numbers $i\in[n]$ modulo $n$. We call a context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$, where $n\geq 3$, a chordless-cycle context set if for all $i,j\in[n]$, the contexts $C_{i}$ and $C_{j}$ intersect if and only if $i=j\pm 1$. For the remainder of this section, if the context $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$ is implicitly understood, we write $\mathcal{S}_{i}$ as a shorthand for $\mathcal{S}_{C_{i}}$, given a contextual $K$-family $\mathcal{S}$ over $\mathcal{C}$.
Let $\mathcal{R}$ be a contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. The overlap projection graph of $\mathcal{R}$, denoted $\mathrm{OPGraph}(\mathcal{R})$, is the digraph s.t.:
  • The vertices consists of the restrictions $s\!\restriction\!(C_{i}\cap C_{i+1})$, for each $s\in\mathcal{R}_{i}$, $i\in[n]$, and
  • There is a directed edge from $s\!\restriction\!(C_{i}\cap C_{i-1})$ to $s\!\restriction\!(C_{i}\cap C_{i+1})$, for each $s\in\mathcal{R}_{i}$, $i\in[n]$.
[…]
A digraph $G=(V,E)$ is called cyclically $n$-partite if there exists a partition $\{V_{0},\dots,V_{n-1}\}$ of $V$ such that for every edge $(u,v)\in E$ we have $u\in V_{i}$ and $v\in V_{i+1}$. The following lemma is a simple observation.
Lemma 6.
Let $\mathcal{R}$ be a contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. Then $\mathrm{OPGraph}(\mathcal{R})$ is cyclically $n$-partite.
A digraph $G$ has an edge cycle cover if every edge $(u,v)\in E$ belongs to some cycle. We will characterise $K$-realisable contextual families with respect to chordless cycles using this notion. The next example showcases the concepts introduced so far.
Example 7.
Consider again the contextual family in Figure 1. The set $\mathcal{C}=\{C_{1},C_{2},C_{3}\}$, where $C_{1}=\{\texttt{Student},\texttt{Teacher}\}$, $C_{2}=\{\texttt{Teacher},\texttt{Course}\}$, and $C_{3}=\{\texttt{Course},\texttt{Student}\}$, is a chordless-cycle context set. The corresponding overlap projection graph, depicted in Figure 2, has $6$ vertices (one for each value) and $7$ edges (one for each assignment). In the figure, we represent a vertex, that is, a restriction $s\upharpoonright\{x\}$, simply as the unique value $s(x)$ it corresponds to. We note that the graph is cyclically $3$-partite, as we can partition the vertices to students, teachers, and courses with edges between them. Moreover, the vertices Alice, Charlie, and Math form a cycle, and similarly Bob, David, and CS form a cycle. Each edge belongs to some cycle, except for the edge from CS to Alice. The graph thus has no edge cycle cover.
MathAliceCharlieCSBobDavid
Figure 2: Example of an Overlap Projection Graph
To characterise $K$-realisable contextual families over chordless cycles, we need the following lemma, which states that every simple cycle gives rise to a uniformly weighted $K$-relation. Recall that a cycle $(v_{1},\dots,v_{\ell},v_{1})$ in a graph is called simple if does not visit any vertex twice. Let us then call a contextual family $\mathcal{R}$ over a chordless-cycle context set simply cyclic if $\mathrm{OPGraph}(\mathcal{R})$ contains a simple cycle
$v_{1}\xrightarrow{s_{1}}v_{2}\xrightarrow{s_{2}}\dots\xrightarrow{s_{\ell-2}}v_{\ell-1}\xrightarrow{s_{\ell-1}}v_{\ell}\xrightarrow{s_{\ell}}v_{1},$ (1)
where =$\bigcup{\mathcal{R}}=\{s_{1},\dots,s_{\ell}\}$. The next lemma states that any simply cyclic family can be realised as a $K$-family using uniform multiplicities. {restatable}lemmahelp Let $\mathcal{R}$ be a simply cyclic contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. Let $K$ be a positive commutative monoid and $0\neq w\in K$. Then $w\mathcal{R}$ is a contextual $K$-family, and $\mathrm{Supp}(w\mathcal{R})=\mathcal{R}$.
Proof.
We need to prove that $w\mathcal{R}$ is locally consistent. Suppose the simple cycle has length $\ell$, as in (1). Since $\mathrm{OPGraph}(\mathcal{R})$ is cyclically $n$-partite (Lemma 6), $\ell$ must be a multiple of $n$; that is, $n\cdot m=\ell$ for some positive integer $m$. Fix $i\in[n]$. One observes that
$|\mathcal{R}_{i}\!\restriction\!(C_{i-1}\cap C_{i})|=|\mathcal{R}_{i}|=|\mathcal{R}_{i}\!\restriction\!(C_{i}\cap C_{i+1})|=m.$
Moreover, each element from $\mathcal{R}_{i}\!\restriction\!(C_{i}\cap C_{i+1})$ appears in $\mathcal{R}_{i+1}\!\restriction\!(C_{i}\cap C_{i+1})$, and conversely each element of $\mathcal{R}_{i+1}\!\restriction\!(C_{i}\cap C_{i+1})$ appears in $\mathcal{R}_{i}\!\restriction\!(C_{i}\cap C_{i+1})$. We conclude that $(w\mathcal{R})_{i}\!\restriction\!(C_{i}\cap C_{i+1})=(w\mathcal{R})_{i+1}\!\restriction\!(C_{i}\cap C_{i+1})$. If $i$ and $j$ are not adjacent, then $C_{i}\cap C_{j}=\emptyset$. For this case it suffices to note that $(w\mathcal{R})_{i}\!\restriction\!\emptyset$ consists of the empty assignment associated with the sum ${w+\dots+w}$ of length $m$. Thus $w\mathcal{R}$ is a contextual $K$-family. It follows by positivity of $K$ that $\mathrm{Supp}(w\mathcal{R})=\mathcal{R}$. ∎
Theorem 8.
Let $K$ be a positive commutative cancellative monoid and $\mathcal{R}$ a contextual family over a chordless-cycle context set $\mathcal{C}=\{C_{0},\dots,C_{n-1}\}$. Then, $\mathcal{R}$ is the support of some contextual $K$-family if and only if its overlap projection graph has an edge cycle cover.
Proof.
( $\Rightarrow$ ) Let $\mathcal{R}$ be the support of some contextual $K$-family $\mathcal{S}$. Then $\mathcal{S}$ can be represented as #$\#\colon\bigcup{\mathcal{R}}\to K_{\neq 0}$, a mapping which we extend to sets $S$ by #$\#(S)\coloneqq\sum_{s\in S}\#(s)$. Toward a contradiction, assume there is no edge cycle cover for $\mathrm{OPGraph}(\mathcal{R})$. Then, some $s\in\bigcup{\mathcal{R}}$ gives rise to an edge that is not in any cycle. Without loss of generality, we assume that $s$ is associated with the context $C_{0}$, that is, it belongs to $\mathcal{R}_{0}$. We define sets of assignments $L^{j}$, $j\geq 0$, inductively as follows:
  • $L_{0}\coloneqq\{s^{\prime}\in\mathcal{R}_{{1}}\mid{s}\!\restriction\!{(C_{0}\cap C_{1})}={s^{\prime}}\!\restriction\!{(C_{0}\cap C_{1})}\}$,
  • $L^{j+1}$ extends $L^{j}$ with assignments $s^{\prime}\in\mathcal{R}_{{j+1}}$, $j\in[n]$, such that there is $s\in L_{i}\cap\mathcal{R}_{{j}}$ with
    ${s}\!\restriction\!{(C_{j}\cap C_{j+1})}={s^{\prime}}\!\restriction\!{(C_{j}\cap C_{j+1})}.$
Now, let $M\coloneqq\bigcup_{j=1}^{\infty}L^{j}$. Since $\mathcal{R}$ is finite, there exists a positive integer $k$ such that $M=L^{k}$. Furthermore, we write $M_{i}\coloneqq\mathcal{R}_{i}\cap M$, for $i\in[n]$. That is, $M_{i}$ consists of the assignments of $M$ over the context $C_{i}$.
We claim that the following inequalities hold:
#$\#(M_{0}\cup\{s\})\leq\#(M_{1})\leq\ldots\leq\#(M_{n-1})\leq\#(M_{0}),$ (2)
where $\leq$ refers to the natural (pre)order of $K$. Note that if this claim holds, we reach a contradiction. Indeed, note that $s\notin L_{0}$ by hypothesis, hence the transitivity of $\leq$ implies #$\#(M_{0})+\#(\{s\})=\#(M_{0}\cup\{s\})\leq\#(M_{0})$. Thus, there is $c\in K$ such that #$\#(M_{0})+\#(\{s\})+c=\#(M_{0})$. By cancellativity, #($\#(\{s\})+c=0$, and hence by positivity, #$\#(s)=0$. This contradicts the fact that $s\in\bigcup{\mathcal{R}}$, where ##\## is a function from $\bigcup{\mathcal{R}}$ to $K_{\neq 0}$.
Thus it remains to prove (2). First, define
  • $\mathcal{R}_{i,+}=\{s^{\prime}\in\mathcal{R}_{i}\mid s^{\prime}\!\restriction\!{(C_{i}\cap C_{i+1})}\in M_{i}\!\restriction\!(C_{i}\cap C_{i+1})\}$,
  • $\mathcal{R}_{i+1,-}=\{s^{\prime}\in\mathcal{R}_{i+1}\mid s^{\prime}\in{(C_{i}\cap C_{i+1})}\in M_{i}\!\restriction\!(C_{i}\cap C_{i+1})\}.$
Then we observe that, for $i\in\{1,\dots,n-1\}$,
  1. 1.
    $M_{i}\subseteq\mathcal{R}_{i,+}$,
  2. 2.
    #$\#(\mathcal{R}_{i,+})=\#(\mathcal{R}_{i+1,-})$, and
  3. 3.
    $\mathcal{R}_{i+1,-}=M_{i+1}$.
Item 1 is immediate, item 2 follows by local consistency, and item 3 follows by the construction of $M$. Since $\leq$ is transitive and reflexive, we obtain #$\#(M_{i})\leq\#(M_{i+1})$. Also, for the first inequality in (2), we obtain $M_{0}\cup\{s\}\subseteq\mathcal{R}^{\prime}_{0,+}$, #$\#(\mathcal{R}^{\prime}_{0,+})=\#(\mathcal{R}^{\prime}_{1,-})$, $\mathcal{R}^{\prime}_{1,-}=M_{1}$, where $\mathcal{R}^{\prime}_{0,+},\mathcal{R}^{\prime}_{1,-}$ are defined otherwise as $\mathcal{R}_{0,+},\mathcal{R}_{1,-}$, except that $M_{0}$ is replaced with $M_{0}\cup\{s\}$. This concludes the proof of the $\Rightarrow$” direction
( $\Leftarrow$ ) Suppose $\mathrm{OPGraph}(\mathcal{R})$ has an edge cycle cover. Then, each assignment $s\in\bigcup{\mathrm{Supp}(\mathcal{R})}$ belongs to a cycle in $\mathrm{OPGraph}(\mathcal{R})$. In particular, there must be a simply cyclic contextual family $\mathcal{T}_{s}$ such that $s\in\bigcup{\mathrm{Supp}(\mathcal{T}_{s})}\subseteq\bigcup{\mathrm{Supp}(\mathcal{R})}$. Then, Equation 1 entails $w\mathcal{T}_{s}$ is a contextual $K$-family, for any $w\neq 0$ (which exists since $K$ is non-trivial). Since contextual $K$-families over $\mathcal{C}$ are closed under addition (Prop. 3), the sum $\sum_{s\in\bigcup{\mathrm{Supp}(\mathcal{R})}}w\mathcal{T}_{s}$ is a contextual $K$-family. Its support is $\mathcal{R}$, for $K$ is positive. Hence the $\Leftarrow$” follows. ∎
Consider our running example in light of this theorem:
Example 9.
[…]
3.2 Contextual $\mathbb{R}_{\geq 0}$-realisability
Let us next briefly consider the special case of $\mathbb{R}_{\geq 0}$-realisability. As a consequence of the previous section, we observe that a contextual $\mathbb{R}_{\geq 0}$-family decomposes to its cycles. Moreover, we establish that a contextual family is the support of some contextual $\mathbb{R}_{\geq 0}$-family if and only if it is the support of some contextual $\mathbb{N}$-family.
Corollary 10.
Each contextual $\mathbb{R}_{\geq 0}$-family over a chordless-cycle context set is a non-negative combination of its cycles.
Proof.
Let $\mathcal{R}$ be a contextual $\mathbb{R}_{\geq 0}$-family, with a multiplicity mapping #$\#\colon\bigcup{\mathrm{Supp}(\mathcal{R})}\to\mathbb{R}_{>0}$. Assume $\mathcal{R}$ is non-empty; otherwise the claim holds trivially. Theorem 8 entails that there is a non-empty set of assignments $S\subseteq\bigcup{\mathrm{Supp}(\mathcal{R})}$ forming a simple cycle in $\mathrm{OPGraph}(\mathcal{R})$. Let $w=\min\{\#(s)\mid s\in S\}$. Define the updated family $\mathcal{R}^{\prime}\coloneqq\mathcal{R}-wS$, where subtraction is defined in the obvious way. Then, $\mathcal{R}^{\prime}$ is a contextual $\mathbb{R}_{\geq 0}$-family. We repeat this process recursively, setting $\mathcal{R}\coloneqq\mathcal{R}-wS$ at each step, until the family becomes empty. Since $S$ is removed entirely at each step (i.e., it no longer appears in the support), no simple cycle $S$ is used more than once, meaning that the process will terminate. This leads to a decomposition of the form $\mathcal{R}=w_{1}S_{1}+\dots+w_{\ell}S_{\ell}$. ∎
This concludes our discussion of context sets that form chordless cycles. A natural follow-up question is whether a similar graph-theoretic characterisation exists for $\mathbb{R}_{\geq 0}$-realisability over arbitrary context sets. The next example suggests that the answer is negative. We now write $x_{1}\dots x_{n}\mapsto a_{1}\dots a_{n}$ for the assignment that maps $x_{1},\dots,x_{n}$ respectively to $a_{1},\dots,a_{n}$.
Example 11.
Consider the contextual family represented in Figure 3. We note that restricted to $\{ab,bc,ca\}$ the corresponding overlap projection graph has an edge cover, and thus the restriction is $\mathbb{R}_{\geq 0}$-realisable by Theorem 8. Similar reasoning applies to the restriction on $\{ab^{\prime},b^{\prime}c,ca\}$. However, over the full context set, $\mathbb{R}_{\geq 0}$-realisability does not hold anymore. Assume toward a contradiction that this is not the case. Let $S$ be the set of assignments represented in Figure 3. Then, Proposition 12 yields a mapping $\pi\colon S\to\mathbb{N}$ which forms a contextual $\mathbb{N}$-relation and is such that $\pi(bc\mapsto 01)=m\geq 1$. By local consistency, $\pi(ca\mapsto 10)=n\geq m$, and furthermore $\pi(ca\mapsto 10)=\pi(ab^{\prime}\mapsto 00)=\pi(b^{\prime}c\mapsto 00)=\pi(ca\mapsto 01)=\pi(ab\mapsto 11)=\pi(bc\mapsto 11)$. Thus
$\pi(bc\mapsto 01)+\pi(bc\mapsto 11)=n+m\neq n=\pi(ca\mapsto 10),$
violating local consistency.
$\begin{array}[t]{cccccc}\begin{array}[t]{cc}a&b\\ \hline\cr\hline\cr 0&0\\ 1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}b&c\\ \hline\cr\hline\cr 0&0\\ 0&1\\ 1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}c&a\\ \hline\cr\hline\cr 0&1\\ 1&0\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}a&b^{\prime}\\ \hline\cr\hline\cr 0&0\\ 1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[t]{cc}b^{\prime}&c\\ \hline\cr\hline\cr 0&0\\ 1&1\\ \hline\cr\hline\cr\end{array}\end{array}$
Figure 3: Non-realisable contextual family.
Nevertheless, we can show that $\mathbb{R}_{\geq 0}$-realisability coincides with $\mathbb{N}$-realisability.
Proposition 12.
A contextual family $\mathcal{R}$ is the support of some contextual $\mathbb{R}_{\geq 0}$-family if and only if it is the support of some contextual $\mathbb{N}$-family.
Proof.
It suffices to prove the “only-if” direction. For each contextual family $\mathcal{R}$, there exists a system of linear inequalities $\mathbf{A}\mathbf{x}\geq\mathbf{b}$ (with rational coefficients) that has a (real) solution $\mathbf{x}$ if and only if $\mathcal{R}$ is the support of some contextual $\mathbb{R}_{\geq 0}$-family. Namely, the variables in $\mathbf{x}$ represent the “weights” of the assignments in $\mathcal{R}$, and the inequalities capture the local consistency criterion and the non-negativity of these “weights”. Furthermore, it is well-known that if a system of the described form has a real solution, then it has a rational solution (?). Hence $S$ must be the support of some contextual $\mathbb{Q}_{\geq 0}$-relation $\mathcal{R}$, which can be transformed into a contextual $\mathbb{N}$-relation by multiplying the weight of each assignment by a sufficiently large constant. ∎
4 Entailment of Dependence
Next, we focus on the entailment problem of functional dependence on contextual $K$-families. We first summarise the well-known Armstrong’s axioms for functional dependence in the classical setting (?). We then develop the theory in the contextual setting. In particular, we give a sound and complete axiomatisation for unary dependencies and establish that derivability can be decided in polynomial time.
4.1 Classical setting
A functional dependency (FD) is a statement of the form $x\rightarrow y$, where $x,y\subseteq D$ are finite sets of variables. An FD $x\rightarrow y$ is satisfied in a relation $X$ (written $X\models x\rightarrow y$) if for all $s,s^{\prime}\in X$: $s(x)=s^{\prime}(x)$ entails $s(y)=s^{\prime}(y)$. For a set $\Sigma$ of FDs, we write $\Sigma\models x\rightarrow y$ if for all relations $X$: if $X\models\phi$ for all $\phi\in\Sigma$, then $X\models x\rightarrow y$.
? (?) established that the entailment $\Sigma\models x\rightarrow y$ is completely axiomatised by the following set of three axioms, termed Armstrong’s axioms:
Reflexivity. $x\rightarrow y,\quad\text{for }y\subseteq x.$ Augmentation. If $x\rightarrow y$, then for any $z$, $xz\rightarrow yz.$ Transitivity. If $x\rightarrow y$ and $y\rightarrow z$, then $x\rightarrow z.$
4.2 FDs in the Contextual Setting
For a $K$-relation $R$ and an FD $x\rightarrow y$, it is natural to set that $R\models x\rightarrow y$, if $\mathrm{Supp}(R)\models x\rightarrow y$.
For a functional dependency $\phi$, we write $\mathrm{Vars}(\phi)$ for the set of variables occurring in $\phi$. Given a contextual $K$-family $\mathcal{R}$ over a context set $\mathcal{C}$ and an FD $\phi$ such that $\mathrm{Vars}(\phi)\in\mathcal{C}$, we say that $\mathcal{R}$ satisfies $\phi$, written $\mathcal{R}\models\phi$, if $\mathcal{R}_{\mathrm{Vars}(\phi)}\models\phi$. Note that local consistency guarantees that $\mathcal{R}\models\phi$ iff $\mathcal{R}_{C}\models\phi$, for any (all, resp.) context $C$ such that $\mathrm{Vars}(\phi)\subseteq C$. A finite set of FDs $\Sigma$ entails an individual FD $\phi$ over $K$ (written $\Sigma\models_{K}\phi$) if for every contextual $K$-family $\mathcal{R}$ over $\mathcal{C}$, such that $\mathrm{Vars}(\theta)\in\mathcal{C}$ for all $\theta\in\Sigma\cup\{\phi\}$, $\mathcal{R}\models\Sigma$ implies $\mathcal{R}\models\phi$. If $K=\mathbb{B}$, we sometimes write $\Sigma\models\phi$ and say that $\Sigma$ entails $\phi$ (without explicitly mentioning $\mathbb{B}$).
[…]
Contextual transitivity. If $x\rightarrow y$, $y\rightarrow z$ and $xyz\rightarrow xyz$, then $x\rightarrow z.$
Whether there exist sound and complete axioms, or whether (or when) $\mathbb{B}$-entailment, $\mathbb{R}_{\geq 0}$-entailment and $K$-entailment of FDs generally coincide, are left as open questions. We can however note the following simple property, which follows since every support of a contextual $K$-family is a contextual family, and since $\mathcal{R}\models\phi\Leftrightarrow\mathrm{Supp}(\mathcal{R})\models\phi$ holds for all contextual $K$-families $\mathcal{R}$ and FDs $\phi$.
Proposition 13.
$\Sigma\models\phi$ implies $\Sigma\models_{K}\phi$, for any set of FDs $\Sigma\cup\{\phi\}$ and positive commutative monoid $K$.
4.3 Unary FDs with Binary Contexts
Leaving aside general FDs, let us restrict attention to unary FDs. An FD $x\rightarrow y$ is called unary if both $x$ and $y$ are individual variables. Additionally, FDs of the form $x\rightarrow x$ are called context dependencies or CDs. A CD $x\rightarrow x$, where $x$ consists of two variables, is said to be binary. A CD $x\rightarrow x$ is thus a kind of a dummy FD which only states that $x$ is a part of the context set.
For unary FDs with binary CDs, the converse of Proposition 13 holds. The result follows as a side product of our main technical result: a simple sound and complete axiomatisation for the entailment for unary FDs with binary CDs. We only need to append reflexivity with one extra rule, stating that every cycle of unary functional dependencies can be inverted. This property, termed the cycle rule, is presented below. We write $x_{1}\to x_{2}\to\dots\to x_{n-1}\to x_{n}$ as a shorthand for $x_{1}\rightarrow x_{2},x_{2}\rightarrow x_{3},\dots,x_{n-1}\rightarrow x_{n}$.
Cycle rule. For $k\geq 1$, if $x_{1}\to x_{2}\to\dots\to x_{k}\to x_{1},$ then $x_{1}\rightarrow x_{k}$.
Its soundness proof relies on the assumption of local consistency, as we shall shortly see.
We note that the notion of a formal proof has a subtle feature; the intermediate derivations should not introduce new contexts. Otherwise, for instance, it would be possible to derive $x\rightarrow z$ from $x\rightarrow y$ and $y\rightarrow z$ using reflexivity (to obtain $xyz\rightarrow xyz$) and contextual transitivity (to obtain $x\rightarrow z$). In contrast, the question of logical entailment of $x\rightarrow z$ by $x\rightarrow y$ and $y\rightarrow z$ has been formulated over contextual families having contexts $xy,yz,xz$ but not necessarily the context $xyz$.
Definition 14 (Derivation).
Given a set of inference rules $\mathcal{S}$, and a set of FDs $\Sigma\cup\{\phi\}$, we say that $\phi$ is derivable from $\Sigma$ by $\mathcal{S}$, written $\Sigma\vdash_{\mathcal{S}}\phi$, if there is a finite sequence of FDs $\psi_{0},\dots,\psi_{n-1}$ such that
  1. 1.
    $\psi_{n}=\phi$
  2. 2.
    for $i\in[n]$, $\psi_{i}$ is either from $\Sigma$ or obtained from $\{\psi_{0},\dots,\psi_{i-1}\}$ by using one of the rules of $\mathcal{S}$.
  3. 3.
    for $i\in[n]$, $\mathrm{Vars}(\psi_{i})\subseteq\mathrm{Vars}(\theta)$, for some $\theta\in\Sigma\cup\{\phi\}$.
Specifically, Item 3 is the new criterion introduced for the contextual setting. As usual, we call $\mathcal{S}$ complete if $\Sigma\models\phi$ entails $\Sigma\vdash_{\mathcal{S}}\phi$, and sound if $\Sigma\vdash_{\mathcal{S}}\phi$ entails $\Sigma\models\phi$. By $\mathcal{CR}$ we refer to the set consisting of reflexivity and the cycle rule for each $k\geq 1$.
Lemma 15.
Let $K$ be a positive commutative monoid. The cycle rule is sound for contextual $K$-families, i.e., $x_{1}\to x_{2}\to\dots\to x_{k}\to x_{1}\models_{K}x_{1}\rightarrow x_{k}$, for all $k\geq 1$.
Proof.
We prove the case $K=\mathbb{B}$; the general case then follows from Proposition 13. Assuming the rule is not sound, we construct a contradiction with the assumption that the contextual family is finite. The contradiction is depicted in Figure 4. Let $\mathcal{R}$ be a contextual family that satisfies $x_{1}\to x_{2}\to\dots\to x_{k}\to x_{1}$ but does not satisfy $x_{1}\rightarrow x_{k}$. Now, for a relation over two variables $x_{i},x_{j}$, let us write $(a_{i,\ell},a_{j,\ell^{\prime}})$ to denote the assignment $(x_{i},x_{j})\mapsto(a_{i,\ell},a_{j,\ell^{\prime}})$. By hypothesis, $\mathcal{R}_{x_{1},x_{k}}$ contains two assignments $(a_{1,1},a_{n,1})$ and $(a_{1,1},a_{n,2})$, where $a_{n,1}\neq a_{n,2}$. By local consistency and the FD $x_{k-1}\to x_{k}$, we observe $\mathcal{R}_{x_{k-1},x_{k}}$ then contains two assignments $(a_{n-1,1},a_{n,1})$ and $(a_{n-1,1},a_{n,2})$, where $a_{n-1,1}\neq a_{n-1,2}$. By induction we obtain that $\mathcal{R}_{x_{1},x_{2}}$ contains also two assignments of the form $(b_{1,1},a_{2,1})$ and $(b_{1,2},a_{2,2})$, where $b_{1,1}\neq b_{1,2}$ and $a_{2,1}\neq a_{2,2}$.
Then, choosing $a_{1,2}\in\{b_{1,1},b_{1,2}\}\setminus\{a_{1,1}\}$, local consistency and the FD $x_{k}\rightarrow x_{1}$ entails that $\mathcal{R}_{x_{1},x_{k}}$ contains an assignment of the form $(a_{n,3},a_{1,2})$, where $a_{n,3}\notin\{a_{n,1},a_{n,2}\}$. Now similarly to the above, by induction $\mathcal{R}_{x_{1},x_{2}}$ contains an assignment of the form $(b_{1,3},a_{2,3})$, where $b_{1,3}\notin\{b_{1,1},b_{1,2}\}$. In particular, the elements $b_{1,1},b_{1,2},b_{1,3}$ are distinct. Then, we choose $a_{1,3}\in\{b_{1,1},b_{1,2},b_{1,3}\}\setminus\{a_{1,1},a_{1,2}\}$, and continue as before. Thus, we see that $\mathcal{R}$ must contain relations whose supports are infinite, which contradicts the definition. Therefore we conclude that the rule is sound.
$x_{1}$ $x_{2}$
$b_{1,1}$ $a_{2,1}$
$b_{1,2}$ $a_{2,2}$
[…]
$\dots$
$x_{k-1}$ $x_{k}$
$a_{n-1,1}$ $a_{n,1}$
$a_{n-1,2}$ $a_{n,2}$
[…]
$x_{k}$ $x_{1}$
$a_{n,1}$ $a_{1,1}$
$a_{n,2}$ $a_{1,1}$
[…]
Figure 4: Counterexample for the falsification of the cycle rule.
Notice that the previous proof constructs a counterexample showing that the cycle rule is not sound if relations with infinite support are allowed. Restricting to unary FDs and binary CDs, we next show that the cycle rule together with reflexivity forms an infinite sound and complete proof system.
Theorem 16.
For every positive commutative monoid $K$, $\Sigma\models_{K}\phi$ iff $\Sigma\vdash_{\mathcal{CR}}\phi$, assuming $\Sigma\cup\{\phi\}$ is a set of unary FDs and binary CDs. Specifically, the cycle rule and reflexivity are sound and complete for unary FDs and binary CDs.
Proof.
We prove the case $K=\mathbb{B}$; the proof of the general case is identical. The only change required for the general case is that all relation constructions have to be replaced with uniformly annotated $K$-relations, that is every assignment is annotated with the same monoid value in $K$. Also, note that $\mathcal{R}\models x\rightarrow y$ iff $\mathrm{Supp}(\mathcal{R})\models x\rightarrow y$ holds by definition.
First note that, if $\phi$ is a CD, then both $\Sigma\models\phi$ and $\Sigma\vdash_{\mathcal{CR}}\phi$ trivially hold. Thus we assume that $\phi$ is a unary FD. Let us consider soundness first. Suppose $(\psi_{1},\dots,\psi_{n})$ is a proof of $\psi_{n}$ from $\Sigma$, and let $\mathcal{R}$ be a contextual family over $\mathcal{C}$ such that $\mathcal{R}\models\Sigma$, and $\mathrm{Vars}(\theta)\in\mathcal{C}$ for all $\theta\in\Sigma\cup\{\psi_{n}\}$. We need to show that $\mathcal{R}\models\psi_{n}$. As the induction hypothesis, suppose $\mathcal{R}\models\psi_{i}$ for all $i<j$, where $j\leq n$. Suppose also $\psi_{j}$ is obtained from $\psi_{\ell_{1}},\dots,\psi_{\ell_{n}}$, where $\ell_{1}<\dots<\ell_{n}<j$, by using the cycle rule. In particular, the induction hypothesis entails $\mathcal{R}$ satisfies all of $\psi_{\ell_{1}},\dots,\psi_{\ell_{n}}$ and their contexts are in $\mathcal{C}$ (due to Item 3 of Definition 14). Hence $\mathrm{Vars}(\psi_{j})\in\mathcal{C}$, due to the form of the cycle rule. From Lemma 15 it now follows that $\mathcal{R}\models\psi_{j}$, for $\mathcal{R}$ contains all the necessary contexts. The case for reflexivity is immediate by Item 3 of Definition 14, concluding the induction proof.
For completeness, we prove the contraposition: $\Sigma\not\models x\rightarrow y$ if $\Sigma\not\vdash x\rightarrow y$. Assume $\Sigma\not\vdash x\rightarrow y$. By reflexivity, $x$ and $y$ are different variables. Let $V$ be the set of variables that appear in $\Sigma\cup\{x,y\}$. Consider the directed graph $G$ with node set $V$ and edges consisting of the functional dependencies $u\rightarrow v$ appearing in $\Sigma$. There are two cases:
Case 1. There is no path from $x$ to $y$ in $G$. Let $X$ be the set of variables to which there is a path from $x$, including $x$ itself. We create one relation $R$ satisfying $\Sigma$ and falsifying $x\rightarrow y$. This relation has two assignments: $s$ that maps each variable from $V$ to 00, and $s^{\prime}$ that maps each variable from $X$ to 00 and the remaining variables from $V\setminus X$ to $1$. Since $x\in X$ and $y\in V\setminus X$, we obtain $R\not\models x\rightarrow y$. For every $u\rightarrow v\in\Sigma$, we have $R\models u\rightarrow v$, otherwise $u\in X$ and $v\in V\setminus X$, contradicting the definition of $X$. By taking the projections of $R$ on all subsets of $V$, we obtain a contextual family $\mathcal{R}$ witnessing $\Sigma\not\models x\rightarrow y$.
Case 2. There is a path from $x$ to $y$ in $G$. We construct a counterexample contextual family $\mathcal{R}$ as follows. For each $u\rightarrow v$ or $uv\rightarrow uv$ appearing in $\Sigma$ such that $\{u,v\}\neq\{x,y\}$, construct a relation that has two assignments mapping $(u,v)$ respectively to $(0,0)$ and $(1,1)$. For $x\rightarrow y$, construct a relation that has four assignments mapping $(x,y)$ respectively to $(0,0),(0,1),(1,0),(1,1)$. The local consistency criterion is satisfied, since each variable has values 00 and $1$ in every relation in which it appears. Let $\mathcal{R}$ denote the contextual family that arises from the aforementioned relations. Clearly ⊧̸$\mathcal{R}\not\models x\rightarrow y$, but $\mathcal{R}\models u\rightarrow v$, for each $u\rightarrow v\in\Sigma$ such that $\{u,v\}\neq\{x,y\}$. It remains to consider the case of $u\rightarrow v\in\Sigma$, where $\{u,v\}=\{x,y\}$. Clearly, it cannot be that $u\rightarrow v=x\rightarrow y$, since this would contradict the assumption that $\Sigma\not\vdash x\rightarrow y$. Moreover, if $u\rightarrow v=y\rightarrow x$, then the path from $x$ to $y$ and the cycle rule entails $\Sigma\vdash x\rightarrow y$, contradicting our assumption. Hence we conclude that all FDs in $\Sigma$ are satisfied. The resulting contextual family $\mathcal{R}$ thus provides a proof for $\Sigma\not\models x\rightarrow y$, concluding the second case and the proof. In the general case of $K\neq\mathbb{B}$, using positivity and non-triviality of $K$ we can select two non-zero monoid values $a,b\in K$ such that $a+a=b$. Then assignments in the above defined ($K$-)relations of cardinality $2$ shall be annotated with $b$, while assignments in the ($K$-)relation of cardinality $4$ shall be annotated with $a$. It is easy to see that with this amendment the proof goes through also in the general case. ∎
As the same axioms characterise entailment of unary FDs with respect to both $\mathbb{B}$ and $K$, we obtain the converse of Proposition 13 as a direct corollary.
Corollary 17.
Let $K$ be any positive commutative monoid. Now, $\Sigma\models\phi$ if and only if $\Sigma\models_{K}\phi$, for any set of unary FDs and binary CDs $\Sigma\cup\{\phi\}$.
4.4 Unary FDs with Ternary Contexts
For unary FDs with ternary CDs, the interaction is significantly more intricate. Let us now abbreviate CDs by writing $x$ instead of $x\rightarrow x$. We have already observed that transitivity can fail: from $x_{1}\rightarrow x_{2}$ and $x_{2}\rightarrow x_{3}$, it does not necessarily follow that $x_{1}\rightarrow x_{3}$. However, if we additionally assume $c\rightarrow x_{3}$, $cx_{1}x_{2}$, $cx_{2}x_{3}$, and $cx_{1}x_{3}$, then $x_{1}\rightarrow x_{3}$ does follow. The next rule captures this non-trivial interaction between unary FDs and ternary CDs (the above case is its instance with $n=3$ and $c_{i}=c$ for all $i\in[n]$). Note that even the soundness of the rule is non-trivial.
Contextual chain rule. If 1. $x_{1}\to x_{2}\to\dots\to x_{n-1}\to x_{n},$ 2. $c_{1}\rightarrow x_{n},c_{2}\rightarrow x_{n},\dots,c_{n-1}\rightarrow x_{n}$, 3. $x_{1}c_{1}x_{n}$, 4. $x_{1}c_{1}x_{2},c_{1}x_{2}c_{2},x_{2}c_{2}x_{3},\dots,x_{n-2}c_{n-2}x_{n-1},$
$c_{n-2}x_{n-1}c_{n-1},x_{n-1}c_{n-1}x_{n}$, and
5. $c_{1}c_{2}x_{n},c_{2}c_{3}x_{n},\dots,c_{n-2}c_{n-1}x_{n}$,
then $x_{1}\rightarrow x_{n}$.
Proposition 18.
Let $K$ be a positive commutative monoid. The contextual chain rule is sound for $\models_{K}$.
Proof.
We prove the case $K=\mathbb{B}$; the general case then follows from Prop. 13. Let $\mathcal{R}$ be a contextual family such that the assumptions of the contextual chain rule hold. Suppose toward contradiction that $\mathcal{R}$ does not satisfy $x_{1}\rightarrow x_{n}$. Then $\mathcal{R}$ contains two assignments $s$ and $s^{\prime}$ in context $x_{1}c_{1}x_{n}$ such that $s(x_{1})=s^{\prime}(x_{1})$ and $s(x_{n})\neq s^{\prime}(x_{n})$. Consequently, $c_{1}\rightarrow x_{n}$ entails that $s(c_{1})\neq s^{\prime}(c_{1})$. These two assignments are illustrated in the top-left corner Fig. 5, where without loss of generality we use values 00 and $1$. By local consistency, $\mathcal{R}$ then contains two assignments $t$ and $t^{\prime}$ in the context $x_{1}c_{1}x_{2}$ such that $t(x_{1})=t^{\prime}(x_{1})$ and $t(x_{n})\neq t^{\prime}(x_{n})$. Furthermore, $x_{1}\rightarrow x_{2}$ entails $t(x_{2})=t^{\prime}(x_{2})$. These assignments are illustrated in the top-middle table of Fig. 5, again using values 00 and $1$. By local consistency and induction, the remaining contexts listed in item 4 of the contextual chain rule (ccr) contain assignment pairs of the form depicted in Fig. 5, with the values $b_{i}$ and $b^{\prime}_{i}$ being unknown. In particular, item 1 of ccr guarantees that the two assignments on the remaining relations agree on the variables $x_{i}$, $i\in[n]$.
Moving to the contexts listed in item 5 of ccr, the leftmost table in Fig. 6 illustrates two assignments for the context $c_{1}c_{2}x_{n}$; the values of $c_{1}$ and $c_{2}$ are explained by local consistency with the top-right table of Fig. 5, and the values of $x_{n}$ are explained by $c_{1}\rightarrow x_{n}$ together with local consistency with the top-left table in Fig. 5. The assignments for the context $c_{2}c_{3}x_{n}$ in Fig. 6 arise by local consistency with the assignments for the context $c_{2}x_{2}c_{3}$ in Fig. 5, and from $c_{2}\rightarrow x_{n}$ together with local consistency with the leftmost table in Fig. 6. Using this reasoning, by induction we eventually obtain the rightmost table in Fig. 6.
However, observe now that the pair of variables $(c_{n-1},x_{n})$ is evaluated as $(0,0)$ and $(0,1)$ in Figures 5 and 6, respectively. By local consistency, this leads to a contradiction with the FD $c_{n-1}\rightarrow x_{n}$. We thus conclude by contradiction that $\mathcal{R}$ satisfies $x_{1}\rightarrow x_{n}$.∎
$\begin{array}[]{ccccccc}\begin{array}[]{ccc}x_{1}&c_{1}&x_{n}\\ \hline\cr\hline\cr 0&0&0\\ 0&1&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[]{ccc}x_{1}&c_{1}&x_{2}\\ \hline\cr\hline\cr 0&0&0\\ 0&1&0\\ \hline\cr\hline\cr\end{array}&\begin{array}[]{ccc}c_{1}&x_{2}&c_{2}\\ \hline\cr\hline\cr 0&0&b_{2}\\ 1&0&b^{\prime}_{2}\\ \hline\cr\hline\cr\end{array}\\[21.52771pt] \begin{array}[]{ccc}x_{2}&c_{2}&x_{3}\\ \hline\cr\hline\cr 0&b_{2}&0\\ 0&b^{\prime}_{2}&0\\ \hline\cr\hline\cr\end{array}&\begin{array}[]{ccc}c_{2}&x_{2}&c_{3}\\ \hline\cr\hline\cr b_{2}&0&b_{3}\\ b^{\prime}_{2}&0&b^{\prime}_{3}\\ \hline\cr\hline\cr\end{array}&\dots&\begin{array}[]{ccc}x_{n-1}&c_{n-1}&x_{n}\\ \hline\cr\hline\cr 0&b_{n-1}&0\\ 0&b^{\prime}_{n-1}&0\\ \hline\cr\hline\cr\end{array}\end{array}$
Figure 5: Relations arising from Items 1, 3, and 4.
$\begin{array}[]{ccccccc}\begin{array}[]{ccc}c_{1}&c_{2}&x_{n}\\ \hline\cr\hline\cr 0&b_{2}&0\\ 1&b^{\prime}_{2}&1\\ \hline\cr\hline\cr\end{array}&\begin{array}[]{ccc}c_{2}&c_{3}&x_{n}\\ \hline\cr\hline\cr b_{2}&b_{3}&0\\ b^{\prime}_{2}&b^{\prime}_{3}&1\\ \hline\cr\hline\cr\end{array}&\dots&\begin{array}[]{ccc}c_{n-2}&c_{n-1}&x_{n}\\ \hline\cr\hline\cr b_{n-2}&b_{n-1}&0\\ b^{\prime}_{n-2}&b^{\prime}_{n-1}&1\\ \hline\cr\hline\cr\end{array}\end{array}$
Figure 6: Relations arising from Items 2 and 5.
The proof of the previous proposition suggests that even soundness of potential inference rules can be non-trivial to check for general FDs. We leave it for future work to analyse in more detail entailments between general FDs over contextual families.
We conclude this section with the observation that the contextual transitivity rule is an instance of the contextual chain rule. This implies that reflexivity, augmentation, and contextual chain rule constitute a sound axiomatisation in the contextual setting, that is complete for derivations within a single context. We write $\mathcal{NRA}$ for this set of axioms.
Proposition 19.
Contextual transitivity rule is an instance of the contextual chain rule.
Proof.
It is easy to check that the contextual transitivity rule is obtained from the contextual chain rule by setting $n=3$, $x_{1}=x$, $x_{2}=y$, $x_{3}=z$, $c_{1}=y=c_{2}$. ∎
Corollary 20.
Let $K$ be a positive commutative monoid and $\Sigma\cup\{\phi\}$ a set of FDs and $x$ a set variables that contains all variables that occur in $\Sigma\cup\{\phi\}$. Then $\Sigma\cup\{x\rightarrow x\}\models_{K}\phi$ if and only if $\Sigma\cup\{x\rightarrow x\}\vdash_{\mathcal{NRA}}\phi$.
4.5 Complexity of Derivation
Next, we turn to the complexity of the entailment or derivability problem. In the classical setting, there is a deterministic linear time algorithm to check whether a given set of FDs entails a given FD (?). In this section, we establish a polynomial-time algorithm for deciding whether a given unary FD can be derived from a given set of CDs and unary FDs by the use of reflexivity, cycle rule, and contextual chain rule. It then follows from Theorem 16 that the entailment problem for unary FDs and binary CDs is in polynomial time also in the contextual setting.
Lemma 21.
Let $\Sigma$ be a set of CDs and unary FDs and let $x,y$ be variables. Then the problem, whether $x\rightarrow y$ can be derived from $\Sigma$ by a single application of the contextual chain rule (page 4.4) can be decided in polynomial time with respect to $\lvert\Sigma\rvert$.
Proof.
The problem can essentially be reduced to reachability in an extended dependency graph of $\Sigma$. In order to apply the contextual chain rule for $x$ and $y$, there needs to be a sequence of variables $x_{2},\dots,x_{n-1}$ and a sequence of variables $c_{1},\dots,c_{n-1}$, such that the conditions of the rule are met. This means in particular that there is a “path” of FDs from $x$ to $y$ such that each “edge” $(x_{i},x_{i+1})$ in that path is witnessed by the respective $c_{i},c_{i+1}$ in the sense of item 4 of the contextual chain rule. We construct the graph $(V,E)$, where $V$ contains all pairs of variables in FDs in $\Sigma$ and $E$ contains an edge between $a$ and $b$, if there are some $c_{a},c_{b}$ satisfying the items 2 and 5 of the contextual chain rule, that witness that edge. We then additionally remove edges that do not satisfy item 5 of that rule and add a start node $s$ and a target node $t$, with edges $(s,(x_{1},b))$, for all $b$ and $((x_{n},a),t)$ for all $a$. Then there exists an $s-t$ path in that graph if and only if a single application of the contextual chain rule yields $x_{1}\rightarrow x_{n}$.
This algorithm is developed formally in Algorithm 1. ∎
Input : unary FDs $\mathcal{FD}$, CDs $\mathcal{C}$, $x$, $y$
$V\leftarrow\{(a,c)\mid a,c\in\mathrm{Vars}\}$ ;
  // $\mathcal{O}(n^{2})$
$C_{y}\leftarrow\{C\in\mathcal{C}\mid y\in C\}$ ;
  // $\mathcal{O}(n)$
$W_{y}\leftarrow\{c\in\mathrm{Vars}\mid\exists C\in C_{y}:c\in C,c\rightarrow y\in\mathcal{FD}\}$ ;
  // $\mathcal{O}(n^{2})$
$E_{y}\leftarrow\{((a,c_{1}),(b,c_{2}))\mid{a\rightarrow b\in\mathcal{FD}},\{a,c_{1},b\},\{c_{1},b,c_{2}\}\in\mathcal{C},c_{1},c_{2}\in W_{y}\}\cup\{((a,c),(y,v))\mid{a\rightarrow y\in\mathcal{FD}},\{a,c,y\},\in\mathcal{C},c\in W_{y}\}$ ;
  // $\mathcal{O}(n^{6})$
1 forall $c_{1}\in W_{y}$ do // $\mathcal{O}(n)$
2       if $\{x,c_{1},y\}\notin C_{y}$ then // $\mathcal{O}(n)$
3             forall $(b,c_{2})\in\mathrm{Vars}\times W_{y}$ do // $\mathcal{O}(n^{2})$
                   $E_{y}\leftarrow E_{y}\setminus\{((x,c_{1}),(b,c_{2}))\}$;
                    // $\mathcal{O}(n^{4})$
4
5             end forall
6
7       end if
8
9 end forall
$E_{y}\leftarrow E_{y}\cup\{(s,(x,c))\mid c\in W_{y}\}\cup\{((y,c),t)\mid c\in W_{y}\}$ ;
  // $\mathcal{O}(n^{4})$
10 if there exists an $s-t$ path in $(V,E_{y})$ then // $\mathcal{O}(n^{4})$
       Accept ;
        // $\mathcal{O}(1)$
11
12else
       Reject ;
        // $\mathcal{O}(1)$
13
14 end if
Algorithm 1 Algorithm for the advanced chain rule
Lemma 22.
Let $\Sigma$ be a set of CDs and unary functional dependencies and let $x,y$ be variables. Then the problem, whether $x\rightarrow y$ can be derived from $\Sigma$ by a single application of the cycle rule (see page  4.3) can be decided in polynomial time with respect to $\lvert\Sigma\rvert$.
Proof.
Interpret the unary FDs in $\Sigma$ as a graph, where the variables are the nodes and the dependencies are the edges. Compute the transitive closure of that graph, then check for all incoming edges $(a,y)$ in that transitive closure, whether $y\rightarrow a\in\Sigma$. If $y\rightarrow a\in\Sigma$, then accept, else reject. ∎
Theorem 23.
Let $\Sigma$ be a set of CDs and unary functional dependencies. The set of all unary FDs that can be derived from $\Sigma$ by using the reflexivity rule, the contextual chain rule and the cycle rule can be computed in polynomial time with respect to $\Sigma$.
Proof.
In order to compute the unary derivation closure $\mathrm{DC}$ of $\Sigma$, we first compute all subsets of size $3$ of CDs in $\Sigma$ and add the corresponding CDs to $\Sigma$, so that the contextual chain rule can be applied. They are derivable by using the reflexivity rule and there are only polynomially many such subsets.
We then proceed iteratively: For all pairs of variables $x,y$, such that $x\rightarrow y\notin\mathrm{DC}$, check, whether $\mathrm{DC}$ entails $x\rightarrow y$ by either of the two other rules rules. If so, add $x\rightarrow y$ to $\mathrm{DC}$. Repeat this process, until $\mathrm{DC}$ remains unchanged for all remaining variable pairs.
There are only polynomially many pairs of variables and each rule takes polynomially many steps to verify, so each pass takes polynomial time. Since there are only polynomially many variables, this derivation closure can be at most polynomially large, implying that it can only change in polynomially many passes in this process. Therefore, the entire process takes polynomial time. ∎
Corollary 24.
Given a set $\Sigma\cup\{\phi\}$ of CDs and unary FDs, it can be decided in polynomial time whether $\phi$ can be derived from $\Sigma$ by using the reflexivity rule, the cycle rule and the contextual chain rule
Corollary 25.
Let $K$ be a positive commutative monoid. Given a set $\Sigma\cup\{\phi\}$ of binary CDs and unary FDs, $\Sigma\models_{K}\phi$ can be decided in polynomial time.
5 Conclusion
We considered realisability and logical inference in a theoretical framework for modelling local consistency in the presence of (potential) global inconsistency. The framework is based on locally consistent families of $K$-relations, for positive commutative monoids $K$. We investigated when a given family of locally consistent relations can be realised as locally consistent $K$-families. In particular, we gave a complete characterisation for the case where $K$ is cancellative and $\mathcal{R}$ is a contextual family over chordless-cycle context sets. Moreover, we established a sound and complete axiomatisation for the entailment of unary FDs in the setting of locally consistent $K$-relations. We discovered two weakenings of the transitivity rule—namely, the chain and cycle rules—that are sound in the contextual setting, while the usual transitivity rule of FDs becomes unsound. Finally, we established a polynomial time algorithm for checking whether a given unary FD is derivable from a given set of CDs and unary FDs by using the reflexivity rule, the cycle rule and the contextual chain rule. For unary FDs and binary CDs, this algorithm is a polynomial time decision method for the related entailment problem as well.
We conclude by listing some interesting questions deserving further study:
  • Does there exist a characterisation akin to Theorem 8 for arbitrary context sets? What is the structure of the equivalence relation on monoids that is defined by $K$-realisability of contextual families (cf. Prop. 12)?
  • Is the full (non-unary) entailment problem for FDs invariant of the monoid (cf. Cor. 17)? Is the entailment problem finitely axiomatisable and decidable? For the Boolean monoid, this is a subproblem of the undecidable implication problem for FDs and inclusion dependencies (?).
Acknowledgments
Miika Hannula was partially supported by the ERC grant 101020762. Juha Kontinen was partially funded by grant 359650 of The Research Council of Finland. Timon Barlag, Nina Pardal, and Jonni Virtema were partially supported by the DFG grant VI 1045-1/1. Timon Barlag was partially supported by the DAAD project 57710940.